A sample of glucose, C6H12O6, contains 1.290×1021 carbon atoms.

A.) How many atoms of hydrogen does it contain? Express your answer using four significant figures.

B.) How many molecules of glucose does it contain? Express your answer using four significant figures.

C.) How many moles of glucose does it contain? Express your answer using four significant figures.

D.) What is the mass of this sample in grams? Express your answer using four significant figures.
1
Expert's answer
2019-09-23T03:47:34-0400

A) One molecule of glucose contains two times more atoms of hydrogen than atoms of carbon, then


"N(H) = N(C)\times2=1.290\times10^{21} = 2.580\times10^{21} atoms"


B) One molecule of glucose contains 6 carbon atoms, then number of glucose molecules is 6 times less than number of carbon atoms:

"N(C_6H_{12}O_6)= \frac{N(C)}{6} = \frac{1.290\times10^{21}}{6} =2.150\times10^{20}"

C)

"n=\frac{N}{N_A}"

"n(C_6H_{12}O_6)=\frac{N(C_6H_{12}O_6)}{N_A}=\frac{2.150\times10^{20}}{6.022\times10^{23}}=3.570\times10^{-4} mol"

d)

"m=M\times n"

"M(C_6H_{12}O_6) = 180.16\frac{g}{mol}"

"m(C_6H_{12}O_6) = 180.16\times 3.570\times10^{-4} =6.432\times10^{-2}g"